# In search of the electron-phonon contribution to total energy

### v1

@@ IV What is the phonon contribution to the total energy ?
To identify terms that can be called electron-phonon contributions to the total energy, we first identify the terms that can be called phonon contributions to the total energy. At first sight, the second-order contribution $\mathcal{E}^{(2)}_{i}=\mathfrak{E}_{\jmath j}^{[0]}$ is precisely the phonon contribution, solution of the auxiliary harmonic oscillator Hamiltonian for the nuclear dynamics, with the well-known $M_{0}^{-1/2}$ scaling, in which case higher-order terms should not be labeled as phonon contributions.
However, this choice has three drawbacks. The first concerns the electronic inertial mass, especially the one of the core electrons. The core electrons follow rigidly the nuclei movement. Therefore, the mass of core electrons should be added to the bare nuclear mass in Eq. (2). The second drawback comes from the vector potential in Eq. (14). In fact, in the classical equation of motion that stems from the harmonic approximation, a spatially homogeneous Berry curvature simply generates a force-velocity coupling, which does not prevent one to obtain normal modes of vibrations, with well-characterized frequencies [Qin2012, Saito2019, Bistoni2021, Saparov2022, Bonini2023, Ren2024, Royo2025]. The third drawback stems from the fact that the optimal geometry (the average expectation of the nuclear positions for the exact wavefunction) is not exactly the one of the minimum of the BO hypersurface. The BO curvature at the optimal geometry will differ from the one at the BO geometry and would be better suited to define the phonon contribution to the total energy.
@@
is the correction to the inverse electronic mass. Taking into account that this is only the lowest-order correction to the electronic mass, this expression matches the one obtained using the exact factorization technique, see Eqs. (13), (14), (15) and (B10) of Ref. Scherrer2015. We note that Eq. (88) is quadratic in the scaled momenta, and does not depend on the scaled displacements $\boldsymbol{\rho}_{j}$. Such a quadratic contribution can be transferred from $\hat{\mathfrak{H}}^{[1]}_{j\lambda}[\boldsymbol{\rho}_{j}]$ as defined in Eq. (III) or (III) to $\hat{\mathfrak{H}}^{[0]}_{j}[\boldsymbol{\rho}_{j}]$ defined in Eq. (III) without changing the fact that the latter is an harmonic oscillator Hamiltonian. It simply becomes $\lambda$-dependent.
Similarly, quadratic contributions in the scaled displacements $\boldsymbol{\rho}_{j}$ can be transferred from $\hat{\mathfrak{H}}^{[1]}_{j\lambda}[\boldsymbol{\rho}_{j}]$ to $\hat{\mathfrak{H}}^{[0]}_{j}[\boldsymbol{\rho}_{j}]$, as well as mixed contributions that include the scaled momenta and the scaled displacements such as
@@
Moreover, some contributions might be extracted from the terms in $\hat{\mathfrak{H}}^{[1]}_{j\lambda}[\boldsymbol{\rho}_{j}]$ that are higher-order in the combined scaled displacements and scaled momenta, such as $\lambda E^{\rm BO\{\mathbf{3}\}}_{j}\cdot\boldsymbol{\rho}_{j}^{3}$, and might be transferred to $\hat{\mathfrak{H}}^{[0]}_{j}[\boldsymbol{\rho}_{j}]$, in connection to a $\lambda$-dependent change of reference configuration. These are so-called “quasi-harmonic” contributions. Let us consider Eq. (87). Changing the reference configuration from $\{{\mathbf{R}}_{j}^{0}\}$ to some nearby configuration $\{{\mathbf{R}}^{\rm ref}\}$ will increase the BO energy contribution to the total energy, with a dominant effect proportional to the square of the change of configuration. However, such a change of configuration might induce a linear change of $\lambda E^{\rm BO\{\mathbf{3}\}}_{j}\cdot\boldsymbol{\rho}_{j}^{3}|\vartheta_{\jmath j}\rangle_{\boldsymbol{\rho}_{j}}$ that is present twice in Eq. (87). The net result might be a gain in energy. This possibility is especially relevant when one examines the behavior of the optimized volume of the primitive cell. The standard BO approach starts from the volume that minimizes the BO energy. The derivative of the BO energy with respect to a volume change is zero. However, the zero-point energy of phonons creates a modification of this energy, which is incorporated in the total energy. The gradient of the zero-point energy of phonons with respect to a volume change will usually not vanish at the BO optimized configuration. Hence, the re-optimized volume will be linear in the zero-point energy, that is $\mathcal{\lambda}^{2}$. A $\mathcal{\lambda}^{2}$ change in volume will create a $\mathcal{\lambda}^{4}$ change in total energy.
@@ V Remaining contributions to the total energy
The Hamiltonian $\hat{\mathfrak{H}}_{j\lambda}^{[1]}[\boldsymbol{\rho}_{j}]$ includes the Born potential, that is order $\mathcal{O}(\lambda^{2})$ when developed in powers of the scaled displacement:
@@
The displacement-independent contribution to the total energy of Eq. (91) at the lowest order in $\lambda$ is given in Eq. (III) and can be further evaluated since it does not depend on displacement. Using Eqs. (60c), (53), and (17):
$E^{\rm Born}_{j}[\bf{0}]$ $=\langle\vartheta_{\jmath j}|E^{\rm Born}_{j}[\bf{0}]|\vartheta_{\jmath j}\rangle_{\boldsymbol{\rho}_{j}}$
$=\sum_{\kappa\alpha}\bigg\langle\frac{\partial\phi_{j}}{\partial R_{\kappa\alpha}}_{|\bf{0}}\bigg|\frac{\hat{\cal{P}}^{\textrm{BO}}_{\perp j}}{2m_{\kappa}}\bigg|\frac{\partial\phi_{j}}{\partial R_{\kappa\alpha}}_{|\bf{0}}\bigg\rangle.$ (92)
The derivative of the wavefunction with respect to the displacement of a nucleus is a typical quantity present in the first-principles calculations of dynamical matrices and phonons for solids [Gonze1997, Gonze1997b, Gonze2024a]. Similarly, for molecules, the interatomic force constants are found from such calculations, as well as vibrational frequencies. The scalar product of such a quantity with excited electronic states delivers the electron-phonon (or electron-vibration) matrix elements, as shown later. As such, among all the contributions to the total energy that appear in the BO expansion, Eq. (V) is the lowest-order one that can be called an “electron-phonon contribution to the total energy” (or electron-vibration).
@@
that is independent of the vibrational state. The next Born low-order contribution is linear in the scaled displacement, $n=1$ in Eq. (91). Being odd with respect to the inversion of $\boldsymbol{\rho}_{j}$, its expectation value for the vibrational wavefunction $\vartheta_{\jmath j}$ vanishes. It will not contribute to first order in $\Lambda$ but will contribute to second order in $\Lambda$. This contribution might then be called an electron-phonon contribution to total energy as well, albeit of higher order.
Finally, we have seen that the self-energy $\hat{\Sigma}_{j\lambda}(z,[\boldsymbol{\rho}_{j}])$ yields Eq. (88) at lowest-order and corresponds to a mass renormalization of the Hamiltonian. From the expression of $\hat{\cal{T}}^{\textrm{N}}_{\lambda}[\boldsymbol{\rho}_{j}]$, Eq. (54), and its double use in the self-energy, Eq. (62), one sees that first- and second-order derivatives of the BO wavefunctions play an important role in the self-energy, that might also be named an electron-phonon contribution to the total energy. Beyond the lowest-order contribution, Eq. (88), the next order brings one momentum operator multiplied by $\lambda$. Being odd with respect to the inversion of $\boldsymbol{\rho}_{j}$, its expectation value for the dynamical wavefunction $\vartheta_{\jmath j}$ vanishes. The order 6 energy associated with the mass renormalization from Eq. (89) is $\mathcal{E}_{i}^{(6),\rm elm}$ and is given in Eq. (162) of Appendix C.
Altogether, there are five energy terms at order 6, see Appendix C, including an electron-phonon energy contribution, $\mathcal{E}_{i}^{(6),\rm elph}$, a contribution associated with the vector potential (which is zero in materials with time-reversal symmetry), $\mathcal{E}_{i}^{(6),\rm vec}$, an anharmonic energy contribution, $\mathcal{E}_{i}^{(6),\rm anh}$, and a mixed contribution which involves anharmonic and vector potential contributions, as well as one anharmonic and an electron-phonon contribution, $\mathcal{E}_{i}^{(6),\rm mix}$. This concludes our listing of all the energy contributions to the total energy up to order 6 in the mass-scaling and we present a summary of the terms in Table 1.
Table 1: Summary of all the energy terms up to order 6 in the mass scaling, that contribute to the total energy. The main text contains expressions up to order 4, while Appendix C present those of order 6.
@@
$\mathcal{E}_{i}^{(2)}=\mathfrak{E}_{\jmath j}^{[0]}$ (46) $\mathcal{O}(\lambda^{2})$
@@
$\mathcal{E}_{i}^{(6),\rm elph}$ (161) $\mathcal{O}(\lambda^{6})$
$\mathcal{E}_{i}^{(6),\rm elm}$ (162) $\mathcal{O}(\lambda^{6})$
$\mathcal{E}_{i}^{(6),\rm vec}$ (172) $\mathcal{O}(\lambda^{6})$
$\mathcal{E}_{i}^{(6),\rm anh}$ (174) $\mathcal{O}(\lambda^{6})$
$\mathcal{E}_{i}^{(6),\rm mix}$ (173) $\mathcal{O}(\lambda^{6})$
@@ VI Semiconductors and insulators treated within density-functional theory
In Sec. II, we introduced $\phi_{j}\{{\mathbf{R}}\}$ as the many-body electronic for a fixed nuclear configuration, in the Born-Oppenheimer approximation, see Eq. (6). Such a many-body wavefunction can only be computed in the simplest systems. In this section, one approximates it by a Slater determinant formed from the combination of density-functional theory (DFT) one-electron Kohn-Sham wavefunctions. Moreover, the application to gapped periodic solids is considered. As mentioned in the previous section, the obtained results rely on non-degenerate perturbation theory, with an energy separation between the ground-state energy and the first excited-state energy. This justifies our consideration of gapped solids only. Modifications would be needed to deal with metallic periodic solids.
@@
The ground-state Born-Oppenheimer DFT energy for solids includes [Martin2004] a kinetic energy term, a one-electron atomic potential energy term, the Hartree-exchange-correlation (Hxc) energy term, and a nucleus-nucleus interaction energy term, as follows:
@@
where $\hat{V}^{\textrm{e-N}}$ and $\hat{V}^{\textrm{Hxc}}$ are periodic, and the Kohn-Sham wavefunctions are normalized to one in the unit cell. The one-electron nuclear potential can be considered in an all-electron, pseudopotential, or projector-augmented wave framework. In the latter two frameworks, the potential operator has local and non-local contributions. Also, the nucleus-nucleus energy includes an Ewald energy term from periodic positively charged particles placed in a negatively charged homogeneous background, but also possibly a pseudo-core energy [Hamann1979] for the pseudopotential and projector-augmented wave frameworks.
@@
where one has to write expressions for the second-order and fourth-order energies in the case of DFT for a solid. The physical total energy is obtained with $\lambda=\lambda_{0}$. We then perform a transformation to mode coordinate, see Appendix A, and express the second-order contribution to the total energy per unit cell resulting from atomic vibrations (or phonons), Eq. (150), as:
$E^{\textrm{ph}}\equiv\lambda_{0}^{2}\mathcal{E}^{(2)}=\lambda_{0}^{2}\mathfrak{E}_{0}^{\rm ph}=\int_{\rm BZ}\frac{\rm d{\mathbf{q}}}{\Omega^{\rm BZ}}\sum_{\nu}\frac{\omega_{\nu{\mathbf{q}}}}{2},$ (101)
@@
The fourth-order contribution to the total energy due to phonons from Eq. (151) can be written in different ways. Considering normal mode derivatives of the wavefunctions yields
@@
$=$ $\frac{1}{2}\iint_{\rm BZ}\!\frac{\rm d{\mathbf{k}}\rm d{\mathbf{q}}}{(\Omega^{\rm BZ})^{2}}\sum_{n}^{\rm occ}\sum_{\nu}\bigg\langle\frac{\partial\phi_{n{\mathbf{k}}}}{\partial\eta_{\nu{\mathbf{q}}}}\bigg|\frac{\partial\phi_{n{\mathbf{k}}}}{\partial\eta_{\nu{\mathbf{q}}}}\bigg\rangle.$ (102b)
@@
where $\mathbf{R}_{l\kappa}$ is the position vector of the atom $\kappa$ in cell $l$. The phase difference between these two possibilities does not impact observables as they cancel in the scalar product present in Eq. (102a). In order to derive Eq. (102a), the orthogonality of first-order perturbed wavefunctions to the occupied subspace (parallel transport gauge) has been used [Gonze1995a].
@@
$E^{\rm elph}=\frac{1}{2}\iint_{\rm BZ}\frac{\rm d{\mathbf{k}}}{\Omega^{\rm BZ}}\frac{\rm d{\mathbf{q}}}{\Omega^{\rm BZ}}\\ \times\sum_{n}^{\rm occ}\sum_{\kappa\alpha}\frac{1}{M_{\kappa}}\bigg\langle\frac{\partial\phi_{n{\mathbf{k}}}}{\partial R_{\kappa\alpha{\mathbf{q}}}}\bigg|\frac{\partial\phi_{n{\mathbf{k}}}}{\partial R_{\kappa\alpha{\mathbf{q}}}}\bigg\rangle,$ (106)
with a possible spin doubling in the unpolarized spin case.

### v2

@@ IV What is the phonon contribution to the total energy ?
To identify terms that can be called electron-phonon contributions to the total energy, we first identify terms that can be called phonon contributions to the total energy. At first sight, the second-order contribution $\mathcal{E}^{(2)}_{i}=\mathfrak{E}_{\jmath j}^{\textrm{ho}}$ is precisely the phonon contribution, solution of the auxiliary harmonic oscillator Hamiltonian for the nuclear dynamic, with the well-known $M_{0}^{-1/2}$ scaling, in which case higher-order terms should not be labeled as phonon contributions.
However, this choice has three drawbacks. The first concerns the electronic inertial mass, especially the one of the core electrons. The core electrons follow rigidly the nuclei movement. Therefore, the mass of core electrons should be added to the bare nuclear mass in Eq. (2). The second drawback comes from the vector potential in Eq. (16). In fact, in the classical equation of motion that stems from the harmonic approximation, a spatially homogeneous Berry curvature simply generates a force-velocity coupling, which does not prevent one to obtain normal modes of vibrations, with well-characterized frequencies [67, 78, 10, 79, 11, 69, 77]. The third drawback stems from the fact that the optimal geometry (the average expectation of the nuclear positions for the exact wavefunction) is not exactly the one of the minimum of the BO hypersurface. The BO curvature at the optimal geometry will differ from the one at the BO geometry and would be better suited to define the phonon contribution to the total energy.
@@
is the correction to the inverse electronic mass. Taking into account that this is only the lowest-order correction to the electronic mass, this expression matches the one obtained using the exact factorization technique, see Eqs. (13), (14), (15) and (B10) of Ref. 80. We note that Eq. (84) is quadratic in the scaled momenta, and does not depend on the scaled displacements $\boldsymbol{\rho}_{j}$. Such a quadratic contribution can be transferred from $\Delta\hat{\mathfrak{H}}_{j\lambda}[\boldsymbol{\rho}_{j}]$ as defined in Eq. (III) or (III) to $\hat{\mathfrak{H}}^{\textrm{ho}}_{j}[\boldsymbol{\rho}_{j}]$ defined in Eq. (III) without changing the fact that the latter is an harmonic oscillator Hamiltonian. It simply becomes $\lambda$-dependent.
Similarly, quadratic contributions in the scaled displacements $\boldsymbol{\rho}_{j}$ can be transferred from $\Delta\hat{\mathfrak{H}}_{j\lambda}[\boldsymbol{\rho}_{j}]$ to $\hat{\mathfrak{H}}^{\textrm{ho}}_{j}[\boldsymbol{\rho}_{j}]$, as well as mixed contributions that include the scaled momenta and scaled displacements such as
@@
Moreover, some contributions might be extracted from the terms in $\Delta\hat{\mathfrak{H}}_{j\lambda}[\boldsymbol{\rho}_{j}]$ that are higher-order in the combined scaled displacements and scaled momenta, such as $\lambda E^{\rm BO\{\mathbf{3}\}}_{j}\cdot\boldsymbol{\rho}_{j}^{3}$, and might be transferred to $\hat{\mathfrak{H}}^{\textrm{ho}}_{j}[\boldsymbol{\rho}_{j}]$, in connection to a $\lambda$-dependent change of reference configuration. These are so-called “quasi-harmonic” contributions. Let us consider Eq. (79). Changing the reference configuration from $\{{\mathbf{R}}_{j}^{0}\}$ to some nearby configuration $\{{\mathbf{R}}^{\rm ref}\}$ will increase the BO energy contribution to the total energy, with a dominant effect proportional to the square of the change of configuration. However, such a change of configuration might induce a linear change of $\lambda E^{\rm BO\{\mathbf{3}\}}_{j}\cdot\boldsymbol{\rho}_{j}^{3}|\vartheta_{\jmath j}\rangle_{\boldsymbol{\rho}_{j}}$ that is present twice in Eq. (79). The net result might be a gain in energy. This possibility is especially relevant when one examines the behavior of the optimized volume of the primitive cell. The standard BO approach starts from the volume that minimizes the BO energy. The derivative of the BO energy with respect to a volume change is zero. However, the zero-point energy of phonons creates a modification of this energy, which is incorporated in the total energy. The gradient of the zero-point energy of phonons with respect to a volume change will usually not vanish at the BO optimized configuration. Hence, the re-optimized volume will be linear in the zero-point energy, that is $\mathcal{\lambda}^{2}$. A $\mathcal{\lambda}^{2}$ change in volume will create a $\mathcal{\lambda}^{4}$ change in total energy.
@@ V Remaining contributions to the total energy
The Hamiltonian $\Delta\hat{\mathfrak{H}}_{j\lambda}[\boldsymbol{\rho}_{j}]$ includes the Born potential, that is order $\mathcal{O}(\lambda^{2})$ when developed in powers of the scaled displacement:
@@
The displacement-independent contribution to the total energy of Eq. (87) at the lowest order in $\lambda$ is given in Eq. (81) and can be further evaluated since it does not depend on displacement. Using Eqs. (61c), (54), and (19):
$E^{\rm Born}_{j}[\textbf{0}]=\sum_{\kappa\alpha}\bigg\langle\frac{\partial\phi_{j}}{\partial R_{\kappa\alpha}}_{|\bf{0}}\bigg|\frac{\hat{\cal{P}}^{\textrm{BO}}_{\perp j}}{2m_{\kappa}}\bigg|\frac{\partial\phi_{j}}{\partial R_{\kappa\alpha}}_{|\bf{0}}\bigg\rangle.$ (88)
The derivative of the wavefunction with respect to the displacement of a nucleus is a typical quantity present in the first-principles calculations of dynamical matrices and phonons for solids [33, 30, 32]. Similarly, for molecules, the interatomic force constants are found from such calculations, as well as vibrational frequencies. The scalar product of such a quantity with excited electronic states delivers the electron-phonon (or electron-vibration) matrix elements, as shown later. As such, among all the contributions to the total energy that appear in the BO expansion, Eq. (88) is the lowest-order one that can be called an “electron-phonon contribution to the total energy” (or electron-vibration).
@@
that is independent of the vibrational state. The next Born lowest-order contribution is linear in the scaled displacement, $n=1$ in Eq. (87). Being odd with respect to the inversion of $\boldsymbol{\rho}_{j}$, its expectation value for the vibrational wavefunction $\vartheta_{\jmath j}$ vanishes. It will not contribute to first order in $\Delta\hat{\mathfrak{H}}_{j\lambda}[\boldsymbol{\rho}_{j}]$ but will contribute at the next order. This contribution might then be called an electron-phonon contribution to total energy as well, albeit of higher order.
Finally, we have seen that the self-energy $\hat{\Sigma}_{j\lambda}(z,[\boldsymbol{\rho}_{j}])$ yields Eq. (84) at lowest-order and corresponds to a mass renormalization of the Hamiltonian. From the expression of $\hat{\cal{T}}^{\textrm{N}}_{\lambda}[\boldsymbol{\rho}_{j}]$, Eq. (55), and its double use in the self-energy, Eq. (63), one sees that first- and second-order derivatives of the BO wavefunctions play an important role in the self-energy, which might also be named an electron-phonon contribution to the total energy. Beyond the lowest-order contribution, Eq. (84), the next order brings one momentum operator multiplied by $\lambda$. Being odd with respect to the inversion of $\boldsymbol{\rho}_{j}$, its expectation value for the dynamical wavefunction $\vartheta_{\jmath j}$ vanishes. The order 6 energy associated with the mass renormalization from Eq. (85) is $\mathcal{E}_{i}^{(6),\rm elm}$ and is given in Eq. (LABEL:eq:6el) of Sec. LABEL:App:wf_factorizes of the SI.
Altogether, there are five groups of energy terms at order 6, see Sec. LABEL:app:auxH4 of the SI, including an electron-phonon energy contribution, $\mathcal{E}_{i}^{(6),\rm elph}$, terms associated with the vector potential (which is zero in materials with time-reversal symmetry), $\mathcal{E}_{i}^{(6),\rm vec}$, anharmonic energy terms, $\mathcal{E}_{i}^{(6),\rm anh}$, and mixed terms which involve anharmonic and vector potential contributions, as well as one anharmonic and an electron-phonon term, $\mathcal{E}_{i}^{(6),\rm mix}$. This concludes our listing of all the energy contributions to the total energy up to order 6 in the mass scaling and we present a summary of the terms in Table 1. We now turn to the practical evaluation of the Born contribution to the total energy, which we will now refer to as the electron-phonon contribution to the total energy.
Table 1: Summary of all the energy terms up to order 6 in the mass scaling, that contribute to the total energy. The main text contains expressions up to order 4, while Sec. LABEL:app:auxH4 of the SI presents those of order 6.
@@
$\mathcal{E}_{i}^{(2)}=\mathfrak{E}_{\jmath j}^{\textrm{ho}}$ (47) $\mathcal{O}(\lambda^{2})$
@@
$\mathcal{E}_{i}^{(6),\rm elph}$ (LABEL:eq:6elph) $\mathcal{O}(\lambda^{6})$
$\mathcal{E}_{i}^{(6),\rm elm}$ (LABEL:eq:6el) $\mathcal{O}(\lambda^{6})$
$\mathcal{E}_{i}^{(6),\rm vec}$ (LABEL:eq:6vec) $\mathcal{O}(\lambda^{6})$
$\mathcal{E}_{i}^{(6),\rm anh}$ (LABEL:eq:6anh) $\mathcal{O}(\lambda^{6})$
$\mathcal{E}_{i}^{(6),\rm mix}$ (LABEL:eq:6hyb) $\mathcal{O}(\lambda^{6})$
@@ VI Semiconductors and insulators treated within density-functional theory
In Sec. II, we introduced $\phi_{j}\{{\mathbf{R}}\}$ as the many-body electronic wavefunction for a fixed nuclear configuration, in the Born-Oppenheimer approximation, see Eq. (6). Such a many-body wavefunction can only be computed in the simplest systems. In this section, one approximates it by a Slater determinant formed from the combination of density-functional theory (DFT) one-electron Kohn-Sham wavefunctions [50]. Moreover, the application to gapped periodic solids is considered. As mentioned in the previous section, the obtained results rely on non-degenerate perturbation theory, with an energy separation between the ground-state energy and the first excited-state energy. This justifies our consideration of gapped solids only. Modifications would be needed to deal with metallic periodic solids.
@@
The ground-state Born-Oppenheimer DFT energy for solids includes [50] a kinetic energy term, a one-electron atomic potential energy term, the Hartree-exchange-correlation (Hxc) energy term, and a nucleus-nucleus interaction energy term, as follows:
@@
where $\hat{V}^{\textrm{e-N}}$ and $\hat{V}^{\textrm{Hxc}}$ are periodic, and the Kohn-Sham wavefunctions are normalized to one in the unit cell. The one-electron nuclear potential can be considered in an all-electron, pseudopotential, or projector-augmented wave framework. In the latter two frameworks, the potential operator has local and non-local contributions. Also, the nucleus-nucleus energy includes an Ewald energy term from periodic positively charged particles placed in a negatively charged homogeneous background, but also possibly a pseudo-core energy [35] for the pseudopotential and projector-augmented wave frameworks.
@@
where one has to write expressions for the second-order and fourth-order energies in the case of DFT for a solid. The physical total energy is obtained with $\lambda=\lambda_{0}$. We then perform a transformation to normal mode coordinates, see Sec. LABEL:sec:normal_modes of the SI, and express the second-order contribution to the total energy per unit cell resulting from atomic vibrations (or phonons), Eq. (LABEL:eq:Eph), as:
$E^{\textrm{ph}}\equiv\lambda_{0}^{2}\mathcal{E}^{(2)}=\lambda_{0}^{2}\mathfrak{E}_{0}^{\rm ho}=\int_{\rm BZ}\frac{\rm d{\mathbf{q}}}{\Omega^{\rm BZ}}\sum_{\nu}\frac{\omega_{\nu{\mathbf{q}}}}{2},$ (97)
@@
The fourth-order contribution to the total energy due to phonons from Eq. (LABEL:eq:E4elphfinal) can be written in different ways. Considering normal mode derivatives of the wavefunctions yields
@@
$=$ $\frac{1}{2}\iint_{\rm BZ}\!\frac{\rm d{\mathbf{k}}\rm d{\mathbf{q}}}{(\Omega^{\rm BZ})^{2}}\sum_{n}^{\rm occ}\sum_{\nu}\bigg\langle\frac{\partial\phi_{n{\mathbf{k}}}}{\partial\eta_{\nu{\mathbf{q}}}}\bigg|\hat{\mathcal{P}}_{\perp\textrm{occ},\mathbf{k}}^{\rm BO}\bigg|\frac{\partial\phi_{n{\mathbf{k}}}}{\partial\eta_{\nu{\mathbf{q}}}}\bigg\rangle,$ (98b)
with a possible spin doubling in the unpolarized spin case and where $\hat{\mathcal{P}}_{\perp\textrm{occ},\mathbf{k}}^{\rm BO}$ is the projector on the unoccupied states.
@@
where $\mathbf{R}_{l\kappa}$ is the position vector of the atom $\kappa$ in cell $l$. The phase difference between these two possibilities does not impact observables as they cancel in the scalar product present in Eq. (98b).
@@
$E^{\rm elph}=\frac{1}{2}\iint_{\rm BZ}\frac{\rm d{\mathbf{k}}}{\Omega^{\rm BZ}}\frac{\rm d{\mathbf{q}}}{\Omega^{\rm BZ}}\\ \times\sum_{n}^{\rm occ}\sum_{\kappa\alpha}\frac{1}{M_{\kappa}}\bigg\langle\frac{\partial\phi_{n{\mathbf{k}}}}{\partial R_{\kappa\alpha{\mathbf{q}}}}\bigg|\hat{\mathcal{P}}_{\perp\textrm{occ},\mathbf{k}}^{\rm BO}\bigg|\frac{\partial\phi_{n{\mathbf{k}}}}{\partial R_{\kappa\alpha{\mathbf{q}}}}\bigg\rangle.$ (102)